OpenStudy (anonymous):

How do you find the level curves of x + 3y + 2z = 8 given the z values of 0,2,and 4? My book says at z=0, the point is (0,0)...how is that even possible?

OpenStudy (anonymous):

I would reread the book more carefully because (0,0,0) does not satisfy the equation.

OpenStudy (anonymous):

(0,0,4) is a solution.

OpenStudy (anonymous):

oh! so if i tried to graph that on xy plane is that just (0,0)??

OpenStudy (anonymous):

If z=0 then our given equation becomes x + 3y + 0 = 8. What does this look like?

OpenStudy (anonymous):

I mean, does the equation look familiar?

OpenStudy (anonymous):

Maybe the equation would look more familiar in y=mx+b form.

OpenStudy (anonymous):

doesnt x=8, and y = 8/3 then??